# Efficient Long Video Tokenization via Coordinated-based Patch Reconstruction

### v1

@@ 2.1 Encoder
We further encode video features ${\mathbf{e}}$ into factorized triplane representations [4, 64], i.e., ${\mathbf{z}}=[{\mathbf{z}}^{xy},{\mathbf{z}}^{yt},{\mathbf{z}}^{xt}]$, where ${\mathbf{z}}^{xy}$ captures the global content in ${\mathbf{x}}$ across time (e.g., layout and appearance of the scene or object), ${\mathbf{z}}^{yt}$ and ${\mathbf{z}}^{xt}$ capture the underlying motion in ${\mathbf{x}}$ across two spatial axes (see Figure 9 for visualization). This design is efficient because it represents a video with three 2D latent planes instead of 3D latents widely used in prior approaches [9, 61, 52].
@@ 2.2 Decoder
Given the triplane representation ${\mathbf{z}}=[{\mathbf{z}}^{xy},{\mathbf{z}}^{yt},{\mathbf{z}}^{xt}]$, we implement our decoder to learn a mapping from $(x,y,t)$ coordinate to the pixels of corresponding patch. We first divide the video into non-overlapping space-time patches. We note that the configuration of patches, e.g., patch sizes, may differ from the one used in the video encoder. We then randomly sample $(x,y,t)$ coordinates by sampling $N$ patches and extracting the center coordinates of sampled patches. We find that sampling only $3$% of video patches can achieve strong performance (see Table 4 for the effect of sampling).
As inputs to the decoder, we use coordinate-based representations ${\mathbf{h}}$ that are obtained by querying each coordinate from triplane representations via bilinear interpolation. Specifically, let $(i,j,k)$ be one of sampled coordinates. We extract ${\mathbf{h}}^{xy}$ by querying $(i,j)$ from ${\mathbf{z}}^{xy}$, ${\mathbf{h}}^{yt}$ by querying $(j,k)$ from ${\mathbf{z}}^{yt}$, and ${\mathbf{h}}^{xt}$ by querying $(i,k)$ from ${\mathbf{z}}^{xt}$. We concatenate them to get coordinate-based representations ${\mathbf{h}}$.
Given $N$ coordinate-based representations $[{\mathbf{h}}_{1},...,{\mathbf{h}}_{N}]$, our decoder processes them through a series of self-attention layers, enabling each ${\mathbf{h}}_{i}$ to attend to other representations ${\mathbf{h}}_{j}$. This allows the decoder to aggregate and fuse the information from different coordinates. We then use a linear projection layer to process the output from each ${\mathbf{h}}_{i}$ to pixels of the corresponding patch. Finally, we update the parameters of our encoder and decoder to minimize an $\ell_{2}$ loss between the reconstructed pixels and original pixels.
@@ 3 Experiments
  • Can CoordTok efficiently encode long videos? Does encoding long videos lead to efficient video tokenization? (Figures 3, 4 and 1)
  • Can CoordTok learn meaningful tokens that can be used for downstream tasks such as video generation? (Figures 6 and 3.1) Can efficient video tokenization improve video generation models? (Sections 3.1 and 5)
@@ 3.1 Experimental Setup
Table 1: Frame-wise reconstruction quality of image and video tokenizers. We report metrics that measure the quality of reconstructed frames: PSNR, LPIPS, and SSIM, computed using the 128$\times$128 resolution frames reconstructed by image and video tokenizers trained on the UCF-101 dataset [40]. Total # tokens denotes the number of tokens required for encoding 128-frame videos. # Frames denotes number of frames in a video used for training tokenizers. $\downarrow$ and $\uparrow$ denotes whether lower or higher values are better, respectively.
@@
Method Time (s) Mem (GB) TATS [9] 180.7 9.8 LARP [50] 114.3 3.1 PVDM-L [64] 116.9 4.0 HVDM [19] 052.1 3.9 CoordTok-SiT-L/2 (Ours) 009.8 4.5
@@
Evaluation
@@ 3.2 Long video tokenization
We mainly consider tokenizers used in recent image or video generation models as our baselines. We first consider MaskGIT-AE [5], an image tokenizer, as our baseline to evaluate the benefit of using video tokenizers for encoding videos. Moreover, we consider PVDM-AE [64], which encodes a video into factorized triplane representations and decodes all frames at once, as another baseline. Comparison with PVDM-AE enables us to evaluate the benefit of our decoder design because it shares the same latent structure with CoordTok. We further consider recent video tokenizers that encode videos into 3D latents, i.e., TATS-AE [9], MAGVIT-AE-L [61], and LARP [50], as our baselines. For a fair comparison, we train all baselines from scratch on UCF-101 or use the model weights trained on UCF-101 following their official implementations. We provide more details of each baseline in Appendix C.
@@ Results
For qualitative evaluation, we provide videos reconstructed by CoordTok and other baseline tokenizers in Figure 3. We report the rFVD of CoordTok and other tokenizers with varying number of token sizes in Figure 4. Notably, we find that CoordTok efficiently encodes 128-frame videos into only 1280 tokens. In contrast, baselines achieve significantly worse reconstruction quality when they use a similar number of tokens to CoordTok. For instance, CoordTok can encode 128-frame videos to 1280 tokens with a rFVD score of 103, while PVDM-AE achieves $>$1000 rFVD score when using 1152 tokens. This highlights the benefit of our decoder design, which enables the tokenizer to exploit the temporal coherence of long videos better for efficient tokenization. Moreover, Table 1 shows CoordTok outperforms baseline tokenizers across diverse metrics that assess the quality of reconstructed frames.
@@ 3.3 Long video generation
We consider recent video generation models that can generate 128-frame videos as baselines, i.e., MoCoGAN [45], MoCoGAN-HD [44], DIGAN [63], StyleGAN-V [37], PVDM-L [64], and HVDM [19]. We provide more details of each baseline in Appendix C.
Figure 6: Unconditional 128-frame, 128$\times$128 resolution video generation results from CoordTok-SiT-L/2 trained on 128-frame videos from the UCF-101 dataset [40]. We provide more examples of generated videos in Appendix D.
(a) Effect of Model size
(b) Effect of Triplane size (spatial)
(c) Effect of Triplane size (temporal)
@@ Results
Section 3.1 provides the quantitative evaluation of our model, i.e., CoordTok-SiT-L/2, and other video generation models. We find that CoordTok-SiT-L/2 achieves the best FVD score of 369.3, outperforming previous baselines. This is an intriguing result considering that CoordTok-SiT-L/2 can generate 128-frame videos much faster than other baselines, as shown in Section 3.1. Moreover, to investigate whether efficient video tokenization improves video generation, we evaluate the FVD scores of SiT-L/2 models trained with CoordTok using token sizes of 1280 and 3072. Figure 5 shows that SiT-L/2 trained with the token size of 1280 achieves consistently low FVD scores, even though there is no significant difference in the reconstruction quality of CoordTok with 1280 and 3072 tokens (see Appendix E). This is likely because the reduced number of tokens makes it easier to train the SiT model. For qualitative evaluation, we provide videos from CoordTok-SiT-L/2 in Figure 6.
@@ 3.4 Analysis and ablation studies
In Figure 7(b) and Figure 7(c), we investigate the effect of spatial and temporal dimensions in triplane representations. We evaluate CoordTok with varying spatial dimensions, i.e., 16$\times$16, 32$\times$32, and 64$\times$64, and varying temporal dimensions, i.e., 8, 16, and 32. In general, we find that using larger planes improves the quality of reconstructed videos, as the model can better represent details within videos using more tokens. This result suggests there is a trade-off between the number of tokens and the reconstruction quality. In practice, we find that reducing the spatial dimensions to 16$\times$16 while using a high temporal dimension of 32 strikes a good balance, achieving good quality of reconstructed videos with a relatively low number of tokens.
(a) Effect of triplane representations.
(b) Effect of coordinate-based representations
@@ A.1 Long video tokenization
Detailed description of coordinate-based patch reconstruction
We describe the coordinate-based patch reconstruction process of CoordTok during training, which involves: (1) constructing the inputs and target outputs for the decoder, (2) obtaining coordinate-based representations from the inputs, and (3) reconstructing the target outputs using these representations.
  • Decoder input and target. Given a video ${\mathbf{x}}$, we patchify it and set $N$ patch coordinates as inputs and their corresponding patch values as target outputs. Formally, we start by patchifying the video ${\mathbf{x}}$, i.e., $\texttt{Patchify}({\mathbf{x}}):=\{(m,{\mathbf{x}}_{(m)})\}_{m=1}^{M}$. We then convert each patch index $m$ into the coordinates $(i,j,k)$ representing the center position of the patch along each axis relative to the entire video ${\mathbf{x}}$, i.e., $i,j,k\in[0,1]$. Finally, we randomly sample $N$ patches. We note that we use entire $M$ patches for evaluation.
    $&\text{Decoder input:}&&[(i_{1},j_{1},k_{1}),\cdots,(i_{N},j_{N},k_{N})]\\ &\text{Target output:}&&[{\mathbf{x}}_{(i_{1}j_{1}k_{1})},\cdots,{\mathbf{x}}_{(i_{N}j_{N}k_{N})}]\end{aligned}$ (1)
    We also use triplane representation ${\mathbf{z}}=[{\mathbf{z}}^{xy},{\mathbf{z}}^{yt},{\mathbf{z}}^{xt}]$ as input of decoder, where the planes have the shape of $H^{\prime}\times W^{\prime}$, $W^{\prime}\times T^{\prime}$, and $H^{\prime}\times T^{\prime}$, respectively.
  • Coordinate-based representations. We use coordinate-based representations $[{\mathbf{h}}_{1},\cdots,{\mathbf{h}}_{N}]$ obtained by triplane representation as input of the transformer decoder. Formally, let $(i,j,k)\in[0,1]^{3}$ be one of the sampled input coordinates. We extract ${\mathbf{h}}^{xy}$ by querying $(i,j)$ from ${\mathbf{z}}^{xy}$, ${\mathbf{h}}^{yt}$ by querying $(j,k)$ from ${\mathbf{z}}^{yt}$, and ${\mathbf{h}}^{xt}$ by querying $(i,k)$ from ${\mathbf{z}}^{xt}$ as follows:
    $&{\mathbf{h}}^{xy}=\text{Bilinear}((i,j);{\mathbf{z}}^{xy}_{lm},{\mathbf{z}}^{xy}_{l,m+1},{\mathbf{z}}^{xy}_{l+1,m},{\mathbf{z}}^{xy}_{l+1,m+1})\\ &{\mathbf{h}}^{yt}=\text{Bilinear}((j,k);{\mathbf{z}}^{yt}_{mn},{\mathbf{z}}^{yt}_{m,n+1},{\mathbf{z}}^{yt}_{m+1,n},{\mathbf{z}}^{yt}_{m+1,n+1})\\ &{\mathbf{h}}^{xt}=\text{Bilinear}((i,k);{\mathbf{z}}^{xt}_{ln},{\mathbf{z}}^{xt}_{l,n+1},{\mathbf{z}}^{xt}_{l+1,n},{\mathbf{z}}^{xt}_{l+1,n+1})\end{aligned}$ (2)
    where $(l,m,n)=(\lfloor i\times(H^{\prime}-1)\rfloor,\lfloor j\times(W^{\prime}-1)\rfloor,\lfloor k\times(T^{\prime}-1)\rfloor)$, $({\mathbf{z}}^{xy}_{lm},{\mathbf{z}}^{yt}_{mn},{\mathbf{z}}^{xt}_{ln})$ indicates the latent vector in ${\mathbf{z}}$ at coordinates $(l,m,n)$, and $\text{Bilinear}(\cdot;\cdot)$ indicates the bilinear interpolation operation at the input coordinate between given vectors. We then concatenate them to get coordinate-based representation of $(i,j,k)$, i.e., ${\mathbf{h}}:=\text{Concat}({\mathbf{h}}^{xy},{\mathbf{h}}^{yt},{\mathbf{h}}^{xt})$.
  • Patch reconstruction. Given $N$ coordinate-based representations $[{\mathbf{h}}_{1},\cdots,{\mathbf{h}}_{N}]$, the transformer decoder processes them through self-attention layers to reconstruct the patches $[\hat{{\mathbf{x}}}_{(i_{1}j_{1}k_{1})},\cdots,\hat{{\mathbf{x}}}_{(i_{N}j_{N}k_{N})}]$. Before the transformer layers, we add positional embeddings corresponding to the coordinates $(i_{1},j_{1},k_{1}),\cdots,(i_{N},j_{N},k_{N})$ to each patch. This allows the self-attention layers to effectively distinguish each patch.
@@ B.3 Analysis
  • Dynamics magnitude. To measure how dynamic each video is, we use the pixel value differences between consecutive frames. To be specific, we compute the dynamics magnitude for each pair of consecutive frames, calculate the mean of these values, and then take the logarithm. Here, dynamics magnitude of two frames $f^{1}$ and $f^{2}$ of resolution $H\times W$ can be defined as follows:
@@
  • Frequency magnitude. To measure the frequency magnitude, we use the metric proposed in Yan et al. [58] that utilizes a Sobel edge detection filter. To be specific, to get the frequency magnitude, we apply both horizontal and vertical Sobel filters to each frame to compute the gradient magnitude at each pixel. We then calculate the average of these magnitudes across all pixels.
@@ C.1 Long video reconstruction
Method Input shape Token shape MaskGiT-AE [5] 128$\times$128$\times$3 8$\times$8 TATS-AE [9] 16$\times$128$\times$128$\times$3 4$\times$16$\times$16 MAGVIT-AE-L [61] 16$\times$128$\times$128$\times$3 4$\times$16$\times$16 PVDM-AE [64] 16$\times$128$\times$128$\times$3 (16$\times$16) $\times$ 3 LARP [50] 16$\times$128$\times$128$\times$3 1024

### v4

# Efficient Long Video Tokenization via Coordinate-based Patch Reconstruction

@@ 2.1 Encoder
After that, we encode video features ${\mathbf{e}}$ into factorized triplane representations [4, 66], i.e., ${\mathbf{z}}=[{\mathbf{z}}^{xy},{\mathbf{z}}^{yt},{\mathbf{z}}^{xt}]$, where the planes have the shape of $H^{\prime}\times W^{\prime}$, $W^{\prime}\times T^{\prime}$, and $H^{\prime}\times T^{\prime}$, respectively. Intuitively, ${\mathbf{z}}^{xy}$ captures the global content in ${\mathbf{x}}$ across time (e.g., layout and appearance of the scene or object), ${\mathbf{z}}^{yt}$ and ${\mathbf{z}}^{xt}$ capture the underlying motion in ${\mathbf{x}}$ across two spatial axes (see Figure 8 for visualization). This design is efficient because it represents a video with three 2D latent planes instead of 3D latents widely used in prior approaches [11, 63, 54].
@@ 2.2 Decoder
Given the triplane representation ${\mathbf{z}}=[{\mathbf{z}}^{xy},{\mathbf{z}}^{yt},{\mathbf{z}}^{xt}]$, we implement our decoder to reconstruct partial video during the training stage by learning a mapping from $(i,j,k)$ coordinate to the pixels of the corresponding patch.
Input and target
We use patch coordinates as inputs to the decoder and their corresponding patch RGB values as targets. Specifically, we first divide the video ${\mathbf{x}}$ into non-overlapping space-time patches. We note that the configuration of patches, e.g., patch sizes, may differ from the one used in the video encoder. We then convert each patch index into the $(i,j,k)$ coordinates representing the center position of the patch along each $x$, $y$, and $t$ axis relative to the entire video ${\mathbf{x}}$, where $i,j,k\in[0,1]$. Finally, we randomly sample $N$ patches. We find that sampling only $3$% of video patches can achieve strong performance (see Table 4 for the effect of sampling).
$&\text{Input:}&&[(i_{1},j_{1},k_{1}),\cdots,(i_{N},j_{N},k_{N})]\\ &\text{Target:}&&[{\mathbf{x}}_{i_{1}j_{1}k_{1}},\cdots,{\mathbf{x}}_{i_{N}j_{N}k_{N}}]\end{aligned}$ (1)
Coordinate-based representations
As inputs to the transformer decoder, we use coordinate-based representations ${\mathbf{h}}$ that are obtained by querying each input coordinate from triplane representation via bilinear interpolation. Specifically, let $(i,j,k)$ be one of sampled coordinates. We extract ${\mathbf{h}}^{xy}$ by querying $(i,j)$ from ${\mathbf{z}}^{xy}$, ${\mathbf{h}}^{yt}$ by querying $(j,k)$ from ${\mathbf{z}}^{yt}$, and ${\mathbf{h}}^{xt}$ by querying $(i,k)$ from ${\mathbf{z}}^{xt}$. More specifically, let $(l,m,n)$ be the indices in the triplane representation corresponding to $(i,j,k)$, obtained using the floor function, i.e., $(l,m,n)=(\lfloor iH^{\prime}\rfloor,\lfloor jW^{\prime}\rfloor,\lfloor kT^{\prime}\rfloor)$. Then, coordinate-based representations are computed as follows: {adjustwidth}-0.15cm0pt
$&{\mathbf{h}}^{xy}=\text{Bilerp}((i,j);{\mathbf{z}}^{xy}_{lm},{\mathbf{z}}^{xy}_{l,m+1},{\mathbf{z}}^{xy}_{l+1,m},{\mathbf{z}}^{xy}_{l+1,m+1})\\ &{\mathbf{h}}^{yt}=\text{Bilerp}((j,k);{\mathbf{z}}^{yt}_{mn},{\mathbf{z}}^{yt}_{m,n+1},{\mathbf{z}}^{yt}_{m+1,n},{\mathbf{z}}^{yt}_{m+1,n+1})\\ &{\mathbf{h}}^{xt}=\text{Bilerp}((i,k);{\mathbf{z}}^{xt}_{ln},{\mathbf{z}}^{xt}_{l,n+1},{\mathbf{z}}^{xt}_{l+1,n},{\mathbf{z}}^{xt}_{l+1,n+1})\end{aligned}$ (2)
where $({\mathbf{z}}^{xy}_{lm},{\mathbf{z}}^{yt}_{mn},{\mathbf{z}}^{xt}_{ln})$ indicates the latent vector in ${\mathbf{z}}$ at indices $(l,m,n)$, and $\text{Bilerp}(\cdot;\cdot)$ is the bilinear interpolation operation at the input coordinate between given vectors. We then concatenate them to get the coordinate-based representation of $(i,j,k)$, i.e., ${\mathbf{h}}:=\text{Concat}({\mathbf{h}}^{xy},{\mathbf{h}}^{yt},{\mathbf{h}}^{xt})$.
Patch reconstruction
Given $N$ coordinate-based representations $[{\mathbf{h}}_{1},...,{\mathbf{h}}_{N}]$, our decoder processes them through a series of self-attention layers, enabling each ${\mathbf{h}}_{n}$ to attend to other representations ${\mathbf{h}}_{m}$. This allows the decoder to aggregate and fuse the information from different coordinates. We then use a linear projection layer to process the output from each ${\mathbf{h}}_{n}$ to pixels of the corresponding patch ${\mathbf{x}}_{i_{n}j_{n}k_{n}}$. Finally, we update the parameters of our encoder and decoder to minimize an $\ell_{2}$ loss between the reconstructed pixels and original pixels.
@@ 3 Experiments
  • Can CoordTok efficiently encode long videos? Does encoding long videos lead to efficient video tokenization (Figures 3, 4 and 1)
  • Can CoordTok learn meaningful tokens that can be used for downstream tasks such as video generation? (Section 3.1) Can efficient video tokenization improve video generation models? (Section 3.1)
@@ 3.1 Experimental Setup
Table 1: Reconstruction quality of image and video tokenizers. We report metrics that measure the quality of reconstructed videos: PSNR, LPIPS, SSIM, and rFVD, computed using the 128$\times$128 resolution videos reconstructed by image and video tokenizers evaluated on the UCF-101 dataset [42]. All models except CosmosTokenizer [9] are trained on UCF-101. Total # tokens denotes the number of tokens required for encoding 128-frame videos. # Frames denotes number of frames in a video used for training tokenizers. $\downarrow$ and $\uparrow$ denotes whether lower or higher values are better, respectively.
@@
Method Time (s) Mem (GB) TATS [11] 180.7 9.8 LARP [52] 114.3 3.1 PVDM-L [66] 116.9 4.0 HVDM [21] 052.1 3.9 Latte-L/2 [10] 021.4 3.1 CoordTok-SiT-L/2 (Ours) 009.8 4.5
@@
(a) Effect of Model size
(b) Effect of Triplane size (spatial)
(c) Effect of Triplane size (temporal)
@@ 3.2 Long video tokenization
We mainly consider tokenizers used in recent image or video generation models as our baselines. We first consider MaskGIT-AE [5], an image tokenizer, as our baseline to evaluate the benefit of using video tokenizers for encoding videos. Moreover, we consider PVDM-AE [66], which encodes a video into factorized triplane representations and decodes all frames at once, as another baseline. Comparison with PVDM-AE enables us to evaluate the benefit of our decoder design because it shares the same latent structure with CoordTok. We further consider recent video tokenizers that encode videos into 3D latents, i.e., TATS-AE [11], MAGVIT-AE-L [63], LARP [52], OmniTokenizer-DV [53], and OmniTokenizer-CV [53] as our baselines. For a fair comparison, we train all baselines from scratch on UCF-101 or use the model weights trained on UCF-101 following their official implementations. In addition, we compare CoordTok to CosmosTokenizer-CV [9], a state-of-the-art tokenizer, although it is not a directly comparable baseline because it is trained on a large-scale dataset. We provide more details of each baseline in Appendix C.
@@ Results
For qualitative evaluation, we provide videos reconstructed by CoordTok and other baseline tokenizers in Figure 3. Notably, we find that CoordTok efficiently encodes 128-frame videos into only 1280 tokens. In contrast, baselines achieve significantly worse reconstruction quality when they use a similar number of tokens to CoordTok. For instance, CoordTok can encode 128-frame videos to 1280 tokens with a rFVD score of 103, while PVDM-AE achieves $>$1000 rFVD score when using 1152 tokens. This highlights the benefit of our decoder design, which enables the tokenizer to exploit the temporal coherence of long videos better for efficient tokenization. Moreover, Table 1 shows CoordTok outperforms baseline tokenizers across diverse metrics that assess the quality of reconstructed frames.
@@ 3.3 Long video generation
We consider recent video generation models that can generate 128-frame videos as baselines, i.e., MoCoGAN [47], MoCoGAN-HD [46], DIGAN [65], StyleGAN-V [39], PVDM-L [66], HVDM [21], and Latte-L/2 [10]. We provide more details of each baseline in Appendix C.
@@ Results
Section 3.1 provides the quantitative evaluation of our model, i.e., CoordTok-SiT-L/2, and other video generation models. We find that CoordTok-SiT-L/2 achieves the best FVD score of 369.3, outperforming previous baselines. This is an intriguing result considering that CoordTok-SiT-L/2 can generate 128-frame videos much faster than other baselines, as shown in Section 3.1. Moreover, to investigate whether efficient video tokenization improves video generation, we evaluate the FVD scores of SiT-L/2 models trained with CoordTok using token sizes of 1280 and 3072. Figure 5 shows that SiT-L/2 trained with the token size of 1280 achieves consistently low FVD scores, even though there is no significant difference in the reconstruction quality of CoordTok with 1280 and 3072 tokens (see Appendix D). This is likely because the reduced number of tokens makes it easier to train the SiT model. For qualitative evaluation, we provide videos from CoordTok-SiT-L/2 in Appendix F.
(a) Effect of triplane representations.
(b) Effect of coordinate-based representations
@@ 3.4 Analysis and ablation studies
In Figure 6(b) and Figure 6(c), we investigate the effect of spatial and temporal dimensions in triplane representations. We evaluate CoordTok with varying spatial dimensions (16$\times$16, 32$\times$32, and 64$\times$64), and varying temporal dimensions (8, 16, and 32). In general, we find that using larger planes improves the quality of reconstructed videos, as the model can better represent details within videos using more tokens. This result suggests there is a trade-off between the number of tokens and the reconstruction quality. In practice, we find reducing the spatial dimensions to 16$\times$16 while using a high temporal dimension of 32 strikes a good balance, achieving good quality of reconstructed videos with a relatively low number of tokens.
@@ 5 Conclusion
Acknowledgements
This work was supported by Institute for Information & communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program(KAIST); No.RS-2021-II212068, Artificial Intelligence Innovation Hub) and Samsung Electronics Co., Ltd (IO201211-08107-01). PA holds concurrent appointments as a Professor at UC Berkeley and as an Amazon Scholar. This paper describes work performed at UC Berkeley and is not associated with Amazon. YS is supported in part by Multidisciplinary University Research Initiative (MURI) award by the Army Research Office (ARO) grant No. W911NF-23-1-0277. We thank NVIDIA for providing compute resources through the NVIDIA Academic DGX Grant.
@@ B.3 Analysis
  • Dynamics magnitude To measure how dynamic each video is, we use the pixel value differences between consecutive frames. To be specific, we compute the dynamics magnitude for each pair of consecutive frames, calculate the mean of these values, and then take the logarithm. Here, dynamics magnitude of two frames $f^{1}$ and $f^{2}$ of resolution $H\times W$ can be defined as follows:
@@
  • Frequency magnitude To measure the frequency magnitude, we use the metric proposed in Yan et al. [60] that utilizes a Sobel edge detection filter. To be specific, to get the frequency magnitude, we apply both horizontal and vertical Sobel filters to each frame to compute the gradient magnitude at each pixel. We then calculate the average of these magnitudes across all pixels.
@@ C.1 Long video reconstruction
  • OmniTokenizer-DV [53] introduces image-video joint VQGAN that compresses a 17-frame video clip into 3D discrete tokens with more advanced architecture design (e.g., uses both 2D and 3D patch embedding layers to support both image and video tokenization, uses transformer backbone with causal attention layers).
  • OmniTokenizer-CV [53] uses the same architecture design as OmniTokenizer-DV, but replaces the VQ loss with KL loss so that it compresses a 17-frame video clip into 3D continuous latent vectors.
@@
Method Input shape Token shape MaskGiT-AE [5] 128$\times$128$\times$3 8$\times$8 TATS-AE [11] 16$\times$128$\times$128$\times$3 4$\times$16$\times$16 MAGVIT-AE-L [63] 16$\times$128$\times$128$\times$3 4$\times$16$\times$16 PVDM-AE [66] 16$\times$128$\times$128$\times$3 (16$\times$16) $\times$ 3 LARP [52] 16$\times$128$\times$128$\times$3 1024 OmniTokenizer [53] (1+16)$\times$128$\times$128$\times$3 (1+4)$\times$16$\times$16